[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Aug 23, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Aug 23 19:51:36 UTC 2017


HSU Shift Log for Wed Aug 23 2017 15:51:29 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES16 meso request active through 9Z. Evening or  
midnight shift should renew it (remember we are priority 4 now, not 24).
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170823/bd4c0806/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list